# PlainQAFact: Retrieval-augmented Factual Consistency Evaluation Metric for Biomedical Plain Language Summarization

### v2

@@ 5.2 Explanation-Only Evaluation
As shown in Figure 2, PlainQAFact outperforms all existing automatic metrics across all three criteria. It is 5.7 points higher in Kendall’s $\tau$, 7.5 points higher in Pearson, and 4.0 points higher in AUC-ROC compared with the second-best metric, QAFactEval. These results indicate that PlainQAFact captures the factual consistency of elaborative explanations more effectively than prior automatic metrics. While GPT-4o achieves higher overall scores (Kendall’s $\tau$=54.9, Pearson=60.7, AUC-ROC=88.7), it shows much higher variance (std. 7.7) compared to PlainQAFact (std. 0.1), making it less stable for factual consistency evaluation.
Compared to the results in Table 2, AlignScore is competitive on general PLS tasks but drops on evaluating explanation-only sentences. In contrast, PlainQAFact remains consistently strong when we isolate explanation-only content. PlainQAFact detects added elaborative explanations and checks them with domain knowledge retrieval, which helps when factual consistency depends on facts beyond simple restatement from the source. Since 44% of sentences in our curated benchmark are labeled as elaborative by human annotators, we believe that PlainQAFact is the more suitable and robust factual consistency evaluation metric in this scenario. PlainQAFact also outperforms Llama 3.1 on all three criteria (27.9/28.3/67.0), showing that instruction-tuned LLM judges alone are not enough for reliable evaluation of elaborative content.
Component Method Choice Kendall’s $\tau$ Pearson AUC std.
PlainQAFact 65.7 81.3 96.4 0.2
@@
GPT-4o 62.8 77.1 94.3* 0.1
No (retrieve for all) 61.2 74.2 93.2* 0.7
Answer Extraction Llama 3.1 - - - -
GPT-4o 66.9 81.9 97.2 0.6
PromptRank 66.3 80.2 96.0 -
@@
Summary 49.3 59.6 84.7* 0.5
Table 3: Ablation study of PlainQAFact on PlainFact, analyzing the impact of individual components in PlainQAFact. Abs: abstract; TB: Textbooks; SP: StatPearls. AUC represents AUC-ROC. The first row of each component setting represents our best combination. * indicates an improvement of our metric over prior work with 95% confidence interval. We run each setting for five times and report the standard deviation (std.). Note that only the settings using LLMs may produce fluctuated scores.
@@ 5.3 Ablation Study
PlainQAFact consists of several modules, including a fine-tuned classifier, QA modules, a retrieval function, and an answer overlap evaluation process. To understand the contribution of each component, we conduct an ablation study on PlainFact5.3). Table 3 summarizes the ablation results, where each component of PlainQAFact is individually modified to measure its impact on overall performance.
@@
Fine-tuned Classifier Table 3 shows that removing the classifier and retrieving for every sentence does not improve factual consistency assessment. Performance drops across all criteria (Kendall’s $\tau$ from 65.7 to 61.2; Pearson from 81.3 to 74.2; AUC-ROC from 96.4 to 93.2), and the std. increases to 0.7. This approach also requires notably more computation. Using GPT-4o as a simple classifier provides reasonable results but still underperforms compared to our fine-tuned classifier. These findings support the use of a lightweight fine-tuned classifier to trigger retrieval only when necessary.
@@
Answer Extraction For extracting gold answers from plain language summaries, the LLM-based extractors are better than PromptRank [40]. GPT-4o achieves the best overall scores (Kendall’s $\tau$=66.9, Pearson=81.9, AUC-ROC=97.2), with small gains over our Llama-based method. PromptRank performs better on Kendall’s $\tau$ (increases 0.6), but fails on other two criteria compared to our best combination. Although GPT-4o achieves slightly higher performance, the marginal gain does not justify the increased cost, making Llama 3.1 a more practical and cost-effective choice for large-scale factual consistency evaluation.
Component Method Choice Kendall’s $\tau$ Pearson AUC std.
Source Simplification Abs 63.5 79.0 94.8 0.2
Elaborative Explanation Abs 32.8 38.2 72.8* 0.2
TB 31.2 37.2 70.6* 0.1
SP 33.3 37.8 72.5 1.7
TB + SP 36.7 42.0 75.4 0.1
Abs + TB 40.3 47.0 78.2 0.2
Abs + SP 40.5 47.4 78.3 0.2
Abs + TB + SP 44.2 51.3 81.0 0.1
Full Dataset Abs 62.8 77.1 94.3* 0.1
TB 60.8 73.8 93.0* 0.1
SP 61.3 74.2 93.3* 0.1
TB + SP 63.4 76.8 94.8 0.1
Abs + TB 64.3 79.7 95.4 0.2
Abs + SP 65.0 80.3 95.9 0.2
Abs + TB + SP 65.7 81.3 96.4 0.2
@@
Table 4: Ablation study results on the PlainFact, evaluating how different retrieval sources affect various information types using PlainQAFact. The fine-tuned classifier categorizes input sentences as source simplification or elaborative explanation. Overall, the numbers of factual summaries that only include source simplification and elaborative explanation sentences are 198 and 191 respectively, and 198 and 194 for non-factual summaries. Abs: abstracts; TB: Textbooks; SP: StatPearls. AUC represents AUC-ROC. * indicates an improvement of Abs + TB + SP over other settings with 95% confidence interval. We run each setting for five times and report the standard deviation (std.) in the brackets.
@@
Combined Domain Resources for Retrieval The highest performance is achieved when combining all three sources (Abs + TB + SP). This suggests that comprehensive retrieval, which includes the original source abstracts, improves overall factual consistency assessment. Compared to using only external sources (TB + SP), performance drops by 2.3/4.5/1.6 across the three criteria. Using only abstracts (i.e., no elaborative retrieval) also lowers performance (2.9/4.2/2.1). To further examine the effectiveness of retrieval for source simplification and elaborative explanation, Table 4 breaks down the results as follows:
@@
(2) Elaborative explanation: For summaries containing only explanation sentences, abstracts alone are not sufficient (32.8/38.2/72.8), performing much worse than in simplification cases. Adding abstracts with external sources improves performance, with the best results obtained by combining all three sources (Abs + TB + SP).
(3)Full dataset: Retrieval from StatPearls provides better results than retrieval from Textbooks, highlighting the importance of using high-quality, domain-specific knowledge bases for PLS evaluation. Overall, the best combination across all settings is Abs + TB + SP (65.7/81.3/96.4).
@@
Answer Overlap Evaluation BERTScore outperforms the LERC approach introduced in QAFactEval [21], with gains of 1.2/5.4/0.9 across the three evaluation criteria. This performance gap can be attributed to the fundamental differences between the two metrics. LERC is a learned model trained to predict human factual consistency scores, which may reduce its generalizability to new domains or sentence types. In contrast, BERTScore computes embedding-based semantic similarity, allowing it to capture fine-grained semantic overlap between generated and gold answers. Since PlainQAFact evaluates factual consistency by comparing model-generated answers to those extracted from the source, BERTScore’s sensitivity to semantic alignment makes it a more effective and robust choice for this step of the evaluation.
Sentence-level vs. Summary-level Evaluation To further examine the effect of input granularity in factual consistency evaluation, we deactivate the sentence-splitting function and instead pair each plain language summary with an abstract to evaluate factual consistency at the summary level. Our best sentence-level evaluation achieves an AUC-ROC of 96.4, while the summary-level setting performs worse across all three criteria (Kendall’s $\tau$=49.3, Pearson=59.6, AUC-ROC=84.7; std. 0.5). Evaluating at the summary level introduces broader contextual dependencies, which can omit sentence-specific factual errors or create inaccurate entailments. These results suggest that processing one sentence at a time makes QA-based question generation and answering more focused and reduces noise from unrelated context.
@@
Scientific Abstract: Frequent consumption of excess amounts of sugar-sweetened beverages (SSB) is a risk factor for obesity, type 2 diabetes, cardiovascular disease and dental caries… [52]
@@
Scientific Abstract: Users of the norelgestromin-containing patch reported more breast discomfort, dysmenorrhea, nausea, and vomiting. In the levonorgestrel-containing patch trial, patch users reported less vomiting, headaches, and fatigue…[53]
@@
Scientific Abstract: To assess the effects of environmental interventions (excluding taxation) on the consumption of sugar-sweetened beverages and sugar-sweetened milk, diet-related anthropometric measures and health outcomes, and on any reported unintended consequences or adverse outcomes…[52]
@@
Retrieved Knowledge: Implementation should be accompanied by high-quality evaluations using appropriate study designs, with a particular focus on the long-term effects of approaches suitable for large-scale implementation.
Final Answer: long-term effects
@@ 6 Discussion and Conclusion
Our study advances the assessment of hallucinations in the field of PLS by supplying both the first domain expert-annotated biomedical PLS benchmark PlainFact and a novel retrieval-augmented QA-based factual consistency evaluation metric PlainQAFact. Unlike existing biomedical domain corpora that either lack fine-grained labels for elaborative content [54] or focus solely on text simplification (e.g., FactPICO[9]), PlainFact captures sentence-level distinctions, including simplification versus explanation, functional roles, and explicit alignment to source sentences. This granularity not only enables precise error analysis of hallucinations introduced for clarity, but also provides a reusable framework for researchers in other disciplines (e.g., legal, technical) to replicate our annotation protocol and build high-quality factual consistency datasets tailored to their specialized texts. By offering these data and detailed annotation protocol (A), we anticipate community-driven extensions, such as multilingual adaptations or integration with domain-specific ontologies that will democratize development of trustworthy summarization models.
@@
Building on this resource, PlainQAFact differentiates from prior QA-based metrics (e.g., QAFactEval[21], QuestEval[50]) by first classifying whether a sentence requires external verification given elaborative explanations are often ignored by metrics relying solely on source text, and improve the evaluation efficiency without retrieving for every plain language summary instance. Then, retrieving targeted domain knowledge for each elaborative explanation before posing and answering fact-checking questions. The ablations demonstrate that selective retrieval yields substantial gains in Kendall’s $\tau$, Pearson correlation, and AUC-ROC over both end-to-end LLM judges[41] and alignment-based scorers[19], highlighting that judicious incorporation of external evidence is essential for robust factual consistency assessment. Beyond its empirical strengths, PlainQAFact serves as a plug-and-play evaluation metric for future summarization systems. Developers can leverage our classifier to classify sentences, apply retrieval only where necessary, thus containing computational costs and obtain interpretable question–answer pairs that manifest non-factual hallucinations. More importantly, PlainQAFact is developed with the open-source backbone model (i.e., Llama 3.1), ensuring broad use with transparency, reproducibility, and fewer computational costs. Looking forward, extending our proposed evaluation method to interactive human–AI workflows, where clinicians or subject-matter experts verify QA outputs on-the-fly, could further enhance trust and adoption in high-stakes domains, including medicine, healthcare, or law.
@@
Despite these promising results, we consider the following limitations of this study. (1) Our error analysis reveals that misclassifications, particularly between source simplifications and elaborative explanations, can lead to retrieval mismatches and unanswerable QA prompts. These issues underscore the need for more nuanced classification and question generation modules. Although we provide substitute solutions for classifying input summaries or sentences, our fine-tuned classifier is limited to the biomedical domain and specifically designed for PLS tasks, fine-tuned through PlainFact. We acknowledge that the dataset used for model fine-tuning is insufficient, resulting in only moderate classification accuracy. Given the scarcity of high-quality human-annotated data for classifying source simplification and elaborative explanation information, future efforts are needed on developing domain-specific (e.g., law, healthcare, social science, etc.) classifiers [55] for factual consistency evaluation. (2) While our experiments indicate that our retrieval-augmented evaluation metric can improve factual consistency assessment in most of the settings, especially in elaborative explanation evaluation, the computational time increases compared to NLI-based evaluation metrics and LLM-based evaluators. The trade-off between evaluation precision and efficiency suggests that further optimization will be beneficial. We suggest to balance wisely based on the trade-offs we report in this study regarding the evaluation time and accuracy.

### v3

@@ 5.2 Explanation-Only Evaluation
[…]
Compared to the results in Table 2, AlignScore is competitive on general PLS tasks but drops on evaluating explanation-only sentences. In contrast, PlainQAFact remains consistently strong when we isolate explanation-only content. PlainQAFact detects added elaborative explanations and checks them with domain knowledge retrieval, which helps when factual consistency depends on facts beyond simple restatement from the source. Since 44% of sentences in our curated benchmark are labeled as elaborative by human annotators, we believe that PlainQAFact is the more suitable and robust factual consistency evaluation metric in this scenario.
Component Method Choice $\tau$ Pearson AUC (99% CI) std.
PlainQAFact 65.7 81.3 96.4 [93.6, 98.5] 0.2
@@
GPT-4o 62.8 77.1 94.3* [90.8, 97.1] 0.1
No (retrieve for all) 61.2 74.2 93.2* [89.6, 96.2] 0.7
AE Llama 3.1 - - - -
GPT-4o 66.9 81.9 97.2 [95.2, 98.9] 0.6
PromptRank 66.3 80.2 96.0 [94.2, 98.4] -
@@
Summary 49.3 59.6 84.7* [79.6, 89.5] 0.5
Table 3: Ablation study of PlainQAFact on PlainFact, analyzing the impact of individual components in PlainQAFact. $\tau$: Kendall’s $\tau$; AE: answer extraction; Abs: abstract; TB: Textbooks; SP: StatPearls; AUC: AUC-ROC. The first row of each component setting represents our best combination. * indicates a statistically significant difference compared to PlainQAFact ($p<0.01$). We run each setting for five times and report the standard deviation (std.). Note that only the settings using LLMs may produce fluctuated scores.
@@ 5.3 Ablation Study
PlainQAFact consists of several modules, including a fine-tuned classifier, QA modules, a retrieval function, and an answer overlap evaluation process. To understand the contribution of each component, we conduct an ablation study on PlainFact. Table 3 summarizes the ablation results, where each component of PlainQAFact is individually modified to measure its impact on overall performance.
@@
Fine-tuned Classifier Table 3 shows that removing the classifier and retrieving for every sentence does not improve factual consistency assessment. PlainQAFact statistically significantly outperforms the settings using GPT-4o as the classifier and the setting without a classifier ($p<0.01$). Performance drops across all criteria (Kendall’s $\tau$ from 65.7 to 61.2; Pearson from 81.3 to 74.2; AUC-ROC from 96.4 to 93.2), and the std. increases to 0.7. This approach also requires notably more computation. Using GPT-4o as a simple classifier provides reasonable results but still underperforms compared to our fine-tuned classifier. These findings support the use of a lightweight fine-tuned classifier to trigger retrieval only when necessary.
@@
Answer Extraction For extracting gold answers from plain language summaries, the LLM based extractors perform better than PromptRank [40]. GPT-4o achieves the best overall scores (Kendall’s $\tau$=66.9, Pearson=81.9, AUC-ROC=97.2), with only small gains over our Llama-based method. In PlainFact, PromptRank obtains a slightly lower Pearson correlation than PlainQAFact. However, in the ablation study on the CELLS dataset in Table 10, PlainQAFact significantly outperforms PromptRank in both perturbation settings (see details in  H). Overall, we do not find a statistically significant difference between PlainQAFact and the other answer extraction methods. PlainQAFact is statistically tied with PromptRank (AUC-ROC=$96.0$, $p=0.34$) and with GPT-4o AE (AUC-ROC=$97.2$, $p=1.0$). Although GPT-4o achieves slightly higher performance overall, the marginal gain does not justify the increased cost, making Llama 3.1 a more practical and cost-effective choice for large-scale factual consistency evaluation.
Component Method Choice $\tau$ Pearson AUC (99% CI) std.
Simplification Abs 63.5 79.0 94.8 [91.6, 97.4] 0.2
Explanation Abs 32.8 38.2 72.7* [65.8, 79.2] 0.2
TB 31.2 37.2 74.6* [68.0, 80.9] 0.1
SP 33.3 37.8 72.5* [65.9, 79.1] 1.7
TB + SP 36.7 42.0 75.4* [68.8, 81.7] 0.1
Abs + TB 40.3 47.0 78.1* [71.8, 84.1] 0.2
Abs + SP 40.5 47.4 78.3* [71.7, 84.2] 0.2
Abs + TB + SP 44.2 51.3 81.0 [74.9, 86.7] 0.1
Full Dataset Abs 62.8 77.1 94.3* [90.8, 97.1] 0.1
TB 60.8 73.8 94.0* [90.7, 96.7] 0.1
SP 61.3 74.2 93.3* [89.6, 96.4] 0.1
TB + SP 63.4 76.8 94.8* [91.7, 97.3] 0.1
Abs + TB 64.3 79.7 95.4* [92.2, 97.9] 0.2
Abs + SP 65.0 80.3 95.9 [92.9, 98.2] 0.2
Abs + TB + SP 65.7 81.3 96.4 [93.6, 98.5] 0.2
@@
Table 4: Ablation study results on the PlainFact, evaluating how different retrieval sources affect various information types using PlainQAFact. The fine-tuned classifier categorizes input sentences as source simplification or elaborative explanation. Overall, the numbers of factual summaries that only include source simplification and elaborative explanation sentences are 198 and 191 respectively, and 198 and 194 for non-factual summaries. Simplification: Source Simplification; Explanation: Elaborative Explanation; $\tau$: Kendall’s $\tau$; Abs: abstracts; TB: Textbooks; SP: StatPearls; AUC: AUC-ROC. * indicates a statistically significant difference compared to Abs+TB+SP ($p<0.01$). We run each setting for five times and report the standard deviation (std.) in the brackets.
@@
Combined Domain Resources for Retrieval As shown in Table 3, combining all three evidence sources (Abs+TB+SP) significantly outperforms the other two settings ($p<0.01$). This suggests that comprehensive retrieval, which includes the original source abstracts, improves overall factual consistency assessment. Table 4 shows the retrieval ablation study results in detail. In the full dataset setting, Abs+TB+SP achieves AUC-ROC=$96.4$ and Pearson=$81.3$, outperforming all retrieval variants that use only a subset of the retrieval corpus, including TB+SP (AUC-ROC=$94.8$), TB only ($94.0$), SP only ($93.3$), and Abs+TB ($95.4$). To further examine the effectiveness of retrieval for source simplification and elaborative explanation, Table 4 breaks down the results as follows:
@@
(2) Elaborative explanation: For summaries containing only explanation sentences, abstracts alone are not sufficient, performing much worse than in simplification cases. Adding abstracts with external sources improves performance, with the best results obtained by combining all three sources. Abs+TB+SP outperforms the abstract-only baseline Abs (AUC-ROC=$72.7$) and all partial-view variants (TB+SP, TB, SP, Abs+TB, and Abs+SP). The gain over Abs is substantial, with $\Delta\mathrm{AUC-ROC}=8.3$ on (99% CI [3.7, 13.0]), and similarly large improvements in correlation.
(3)Full dataset: Retrieval from StatPearls provides better results than retrieval from Textbooks, highlighting the importance of using high-quality, domain-specific knowledge bases for PLS evaluation. Overall, the best combination across all settings is Abs+TB+SP (65.7/81.3/96.4). Relative to the abstract-only baseline, Abs (AUC-ROC=$94.3$), the gain in discrimination is $\Delta\mathrm{AUC-ROC}=2.1$ (99% CI [0.5, 4.2], where $\Delta$ is Abs+TB+SP minus Abs).
@@
Answer Overlap Evaluation BERTScore ties with the LERC approach introduced in QAFactEval [21], showing no significant difference in AUC-ROC ($p=0.29$). BERTScore achieves slightly higher nominal values in both correlation metrics. There are fundamental differences between the two metrics. LERC is a learned model trained to predict human factual consistency scores, which may reduce its generalizability to new domains or sentence types. In contrast, BERTScore computes embedding-based semantic similarity, allowing it to capture fine-grained semantic overlap between generated and gold answers. Since PlainQAFact evaluates factual consistency by comparing model-generated answers to those extracted from the source, BERTScore’s sensitivity to semantic alignment makes it a more effective and robust choice for this step of the evaluation.
Sentence-level vs. Summary-level Evaluation To further examine how input granularity affects factual consistency evaluation, we deactivate the sentence-splitting function and instead pair each plain language summary with its abstract to evaluate factual consistency at the summary level. Our best sentence level evaluation achieves an AUC-ROC of $96.4$, which is significantly higher than the summary level setting ($p<0.01$). The summary level setting performs worse across all three criteria (Kendall’s $\tau$=$49.3$, Pearson=$59.6$, AUC-ROC=$84.7$; std. $0.5$). Evaluating at the summary level introduces broader contextual dependencies, which can omit sentence-specific factual errors or create inaccurate entailments. These results suggest that processing one sentence at a time makes QA-based factual consistency evaluation more focused and reduces noise from unrelated context.
@@
Scientific Abstract: Frequent consumption of excess amounts of sugar‐sweetened beverages (SSB) is a risk factor for obesity, type 2 diabetes, cardiovascular disease and dental caries… [54]
@@
Scientific Abstract: Users of the norelgestromin‐containing patch reported more breast discomfort, dysmenorrhea, nausea, and vomiting. In the levonorgestrel‐containing patch trial, patch users reported less vomiting, headaches, and fatigue…[55]
@@
Scientific Abstract: To assess the effects of environmental interventions (excluding taxation) on the consumption of sugar‐sweetened beverages and sugar‐sweetened milk, diet‐related anthropometric measures and health outcomes, and on any reported unintended consequences or adverse outcomes…[54]
@@
Retrieved Knowledge: Implementation should be accompanied by high‐quality evaluations using appropriate study designs, with a particular focus on the long‐term effects of approaches suitable for large‐scale implementation.
Final Answer: long‐term effects
@@ 5.4 Error Analysis on PlainFact
5.5 Error Analysis on FareBio
Type Target PLS Sentence QA Extraction Analysis Rationale
NF This performance was identical in the comparative analysis subgroup and was within the range of physicians at different levels of expertise: 0.86-0.97 and 0. Ref: physicians
Q: This performance was identical in the comparative analysis subgroup and was within the range of what at different levels of expertise?
Pred: physicians
The non-factual error is a value mistake. The target sentence incorrectly truncates the specificity range to “0,” which does not appear in the source and therefore being labeled as non-factual. However, our answer extractor fails to identify the numerical values as gold answer to verify its factual consistency.
NF Imaging also showed chronic compression of the coeliac axis with compensatory hypertrophy of the gastroduodenal artery. Ref: gastroduodenal artery
Q: Imaging also showed chronic compression of the coeliac axis with compensatory hypertrophy of what artery?
Pred: gastroduodenal
The target sentence correctly restates information that is directly present in the source, but it is labeled as non-factual in FareBio.
F In summary, the article discusses the potential impact of COVID-19 on the musculoskeletal system… and emphasizes the need for further research to understand the underlying mechanisms and develop appropriate interventions. Ref: interventions
Q: In summary, the article discusses the potential impact of COVID-19 on the musculoskeletal system, particularly in the context of aging and Long-COVID, and emphasizes the need for further research to understand the underlying mechanisms and develop appropriate what?
Pred: multidisciplinary approach
The predicted answer comes from the original source. This target PLS sentence is classified as a simplification; however, the term “interventions” does not appear in the source.
F Some patients experience persistent symptoms after recovering from the initial infection, a condition known as “Long-COVID.” Ref: infection
Q: Some patients experience persistent symptoms after recovering from what condition known as Long-COVID?
Pred: SARS-CoV-2
In the source, SARS CoV 2 always appears together with the word “infection.” Our QA model captures the relevant answer, but our answer overlap evaluation approach does not align this prediction with the gold answer, which results in a lower PlainQAFact.
Table 6: Error analysis of PlainQAFact on the FareBio dataset. NF/F = non-factual or factual; Ref = gold/extracted answer; Q = generated question; Pred = predicted answer.
As shown in Table 6, we provide several failure cases on the FareBio dataset based on the results in Table 2. We rank the PlainQAFact scores for all summaries and select two examples with the lowest scores among those labeled NF and two examples with the lowest scores among those labeled F. The first NF example in Table 6 shows a true factual error. The target plain language sentence changes the numerical range and introduces the value zero, which does not appear in the source. However, our answer extractor does not reliably extract the exact numbers, so PlainQAFact does not capture the fact errors. The second NF example shows a label noise in FareBio. The target sentence restates information that is clearly present in the source, but the sentence is labeled as NF. Third, we find pipeline errors even when the FareBio label is factual. In the third example, the target sentence correctly summarizes the current knowledge and is classified as simplification by our classifier, but the answer extractor extracts the word “interventions,” which does not appear in the source and thus leads to a low score. In the fourth example, the QA module predicts “SARS CoV 2” as the answer, which is consistent with the “infection” described in the plain language sentence, but the final answer overlap evaluation step cannot match this prediction to the gold answer “infection.” Overall, we conclude that most errors of our PlainQAFact come from the answer extraction and answer overlap evaluation stages in the FareBio dataset, while the dataset contains noisy factuality labels.
To mitigate these errors (§5.4, §5.5), we provide several directions for future work. First, misclassification of plain language sentences by the factuality type classifier can be reduced by training on more annotated examples or use advanced low-resource model tuning strategies for text classification [56]. Second, retrieval errors can be mitigated by conducting precise filtering method to select the most relevant snippets, such as limiting the length of retrieved snippets and chunk size, using a combination of dense and sparse retrieval, and expanding the domain corpus (e.g., PubMed) to cover more specialized topics. Third, to handle unanswerable or vague questions, future work can improve the question filtering stage that removes questions that do not have clear grounding in both source retrieved context, or that are too general to support reliable factual checks using heuristic approaches. Finally, for the answer overlap evaluation module, more advanced entailment-based models can be applied to replace the BERTScore semantic matching to check whether the predicted answer is consistent with the gold answer after simple normalization of entities and synonyms. We report the quantitative error analysis of FareBio in  J.
@@ 6 Discussion and Conclusion
Our study advances the assessment of hallucinations in the field of PLS by supplying both the first domain expert-annotated biomedical PLS benchmark PlainFact and a novel retrieval‐augmented QA‐based factual consistency evaluation metric PlainQAFact. Unlike existing biomedical domain corpora that either lack fine‐grained labels for elaborative content [57] or focus solely on text simplification (e.g., FactPICO[9]), PlainFact captures sentence‐level distinctions, including simplification versus explanation, functional roles, and explicit alignment to source sentences. This granularity not only enables precise error analysis of hallucinations introduced for clarity, but also provides a reusable framework for researchers in other disciplines (e.g., legal, technical) to replicate our annotation protocol and build high‐quality factual consistency datasets tailored to their specialized texts. By offering these data and detailed annotation protocol (A), we anticipate community‐driven extensions, such as multilingual adaptations or integration with domain‐specific ontologies that will democratize development of trustworthy summarization models.
@@
Building on this resource, PlainQAFact differentiates from prior QA‐based metrics (e.g., QAFactEval[21], QuestEval[50]) by first classifying whether a sentence requires external verification given elaborative explanations are often ignored by metrics relying solely on source text, and improve the evaluation accuracy without retrieving for every plain language summary instance. Then, retrieving targeted domain knowledge for each elaborative explanation before posing and answering fact‐checking questions. The ablations demonstrate that selective retrieval yields substantial gains in Kendall’s $\tau$, Pearson correlation, and AUC‐ROC over both end‑to‑end LLM judges[41] and alignment‐based scorers[19], highlighting that judicious incorporation of external evidence is essential for robust factual consistency assessment. Beyond its empirical strengths, PlainQAFact serves as a plug‑and‑play evaluation metric for future summarization systems. Developers can leverage our classifier to classify sentences, apply retrieval only where necessary, thus containing computational costs and obtain interpretable question–answer pairs that manifest non-factual hallucinations. More importantly, PlainQAFact is developed with the open-source backbone model (i.e., Llama 3.1), ensuring broad use with transparency, reproducibility, and lower computational costs. Looking forward, extending our proposed evaluation method to interactive human–AI workflows, where clinicians or subject‐matter experts verify QA outputs on‐the‑fly, could further enhance trust and adoption in high‐stakes domains, including medicine, healthcare, or law.
@@
Despite these promising results, we consider the following limitations of this study. (1) Our error analysis reveals that misclassifications, particularly between source simplifications and elaborative explanations, can lead to retrieval mismatches and unanswerable QA prompts. These issues underscore the need for more nuanced classification and question generation modules. Although we provide substitute solutions for classifying input summaries or sentences, our fine-tuned classifier is limited to the biomedical domain and specifically designed for PLS tasks, fine-tuned through PlainFact. We acknowledge that the dataset used for model fine-tuning is insufficient, resulting in only moderate classification accuracy. Given the scarcity of high-quality human-annotated data for classifying source simplification and elaborative explanation information, future efforts are needed on developing domain-specific (e.g., law, healthcare, social science, etc.) classifiers [56] for factual consistency evaluation. (2) While our experiments indicate that our retrieval-augmented evaluation metric can improve factual consistency assessment in most of the settings, especially in elaborative explanation evaluation, the computational time increases compared to NLI-based evaluation metrics and LLM-based evaluators. The trade-off between evaluation precision and efficiency suggests that further optimization will be beneficial. We suggest to balance wisely based on the trade-offs we report in this study regarding the evaluation time and accuracy. PlainQAFact is a research-oriented metric that prioritizes interpretability and transparency over speed. (3) Given the fair IRA scores, especially for the factuality type field, we acknowledge that our classification method is not fully reliable. The annotated dataset reflects subjective judgments, particularly when distinguishing source simplification from elaborative explanation, and the training set is limited in size. We advise users to apply our pretrained classifier with caution when using it for later factual consistency evaluation steps.